OpenStudy (anonymous):

I need help with he following: Use the given values (if possible) the remaining trig fxns. tan x= 7/24 , sec x= -25/24

OpenStudy (anonymous):

us the given values to evaluate*

OpenStudy (anonymous):

use*

hero (hero):

tan^-1(7/24) sec^-1(-25/24)

OpenStudy (anonymous):

I dont understand how you got that though.

OpenStudy (anonymous):

ok. we can find the hypotenuse from tan x:\[c^2=a^2+b^2=7^2+24^2=625\]So, c=25. Now just do the usual:\[\sin x=\frac{7}{25}\]...etc

OpenStudy (anonymous):

This is about using identities though, i dont see how finding the hypotenuse relates with the such.

hero (hero):